Stock in Thanksgiving Industries has a beta of 1.2. The market risk premium is 8 percent, and T-bills are currently yielding 5.1 percent. The company’s most recent dividend was $1.50 per share, and dividends are expected to grow at an annual rate of 8 percent indefinitely. If the stock sells for $40 per share, what is your best estimate of the company’s cost of equity? Hint: use both approaches! (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

cost of equity as per growth model=(D1/Current price)+Growth rate

=[(1.5*1.08)/40]+0.08=12.05%

cost of equity as per CAPM=risk free rate+Beta*market risk premium

=5.1+(1.2*8)=14.7%

Best estimate of cost of equity=(12.05+14.7)/2

=13.38%(Approx).